OpenStudy (anonymous):

HELP ME PLS:) -(-2x^6y^7)^3 (x^2-y^2)^3

OpenStudy (anonymous):

I'm here again! hahah.

OpenStudy (anonymous):

how great GOD is, hehehe.. so lucky today..

OpenStudy (anonymous):

\[(8x^{18}y^{21}) (x^4 -2x^2y^2 +y^4)(x^2 - y^2) \]

OpenStudy (anonymous):

Hahah. I know, right?

OpenStudy (anonymous):

LOL.

OpenStudy (anonymous):

Then

OpenStudy (anonymous):

then?:)

OpenStudy (anonymous):

\[(8x^{18}y^{21})(x^6 -2x^4y^2 +x^2y^4 -x^4y^2 +2x^2y^4-y^6)\]

OpenStudy (anonymous):

in \[(2x^6-y^7)^3\] i get 8x^18-12x^12y^7+6x^6y^14-y^21 i think im wrong

OpenStudy (anonymous):

then just distribute it. I'm tired of typing. :))

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

\[ 8x^18- 4x^12y^7 +4x^{12}y^7 -2x^6y^{14} + 2x^6y^14-y^21 \]

OpenStudy (anonymous):

my ans. in (x^2-y^2)^3 is \[x^6-3x^4y^2+3x^2y^4-y^6 \]

OpenStudy (anonymous):

I think. LOL

OpenStudy (anonymous):

Oh right right. LOL. Sorry.

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

Yeah. And I'm wrong. XD

OpenStudy (anonymous):

what is your goal? multiply out?

OpenStudy (anonymous):

I gotta go now. For real. hahah.

OpenStudy (anonymous):

satellite, yeah!

OpenStudy (anonymous):

ok first off first term is \[8x^{18}y^{21}\]

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

that is just cubing the first term. then you have to expand the second term. then you have to multiply

OpenStudy (anonymous):

i cant clearly understand:(

OpenStudy (anonymous):

ok you have two things here. one is \[-(-2x^6y^7)^3\] yes?

OpenStudy (anonymous):

and when we cube it we get what i wrote above namely \[8x^{18}y^{21}\]

OpenStudy (anonymous):

multiply the exponents and cube the 2 and there are two minus signs so we remove them

OpenStudy (anonymous):

the next term is \[(x^2-y^2)^3\] and when you cube this thing you get \[x^6-3 x^4 y^2+3 x^2 y^4-y^6\]

OpenStudy (anonymous):

you can multiply twice if you like or you can use \[(a+b)^3=a^3+3a^2b+3ab^3+b^3\] with \[a=x^2,b=-y^2\]

OpenStudy (anonymous):

how are we doing so far?

OpenStudy (anonymous):

it doesnt mean that im going to multiply the terms in itself? for example 1st term... am i not going to multiply by itself thrice? when it doesnt have operations?

OpenStudy (anonymous):

heheheh:) what's the next step?

OpenStudy (anonymous):

not sure exactly what you mean but \[(x^2-y^2)^3=(x^2-y^2)(x^2-y^2)(x^2-y^2)\] and when you multiply out and combine like terms you get \[x^6-3 x^4 y^2+3 x^2 y^4-y^6\]

OpenStudy (anonymous):

last step is to write \[8x^{18}y^{21}(x^6-3 x^4 y^2+3 x^2 y^4-y^6)\] and multiply out using distributive law

OpenStudy (anonymous):

i get that.. now I know.. thank you:) i already forgot my 2nd yr. lesson.. thank you very much:)

OpenStudy (anonymous):

just do the for multiplications adding the exponents as you go. when you are finished you should have \[8 x^{24} y^{21}-24 x^{22} y^{23}+24 x^{20} y^{25}-8 x^{18} y^{27}\]

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

is that when we are adding in algebraic ex. we cannot add the terms if they have different exponents?

OpenStudy (anonymous):

that is the final answer?

OpenStudy (anonymous):

this is the final answer

OpenStudy (anonymous):

yes you cannot add the terms if the have different exponents....the terms are linearly independent